OpenStudy (anonymous):

Given the function g(x) = −3x + 4, compare and contrast g(−2) and g(4). Choose the statement that is true concerning these two values.

jhonyy9 (jhonyy9):

g(-2)=6+4=10 g(4)=-12+4=-8